(Derivation of perpetuity formula) Using the geometric series formula: if −1 < r < 1, then...

  1. (Derivation of perpetuity formula) Using the geometric series formula: if −1 < r < 1, then

    a+ar+ar2 +ar3 +···= a ,1−r

    derive the PV of perpetuity formula.
Answer #1

annual payment = P

discount rate = r

PV = P/(1 + r) + P/(1 + r)2.........

(first term (a) = P/(1 + r) and common ratio (x) = 1/(1 + r))

(sum of infinite GP = a/(1 - x))

PV = P/(1 + r) x (1/(1 - 1/(1 + r))

PV = P/r